# HPC Alongside User-space Kubernetes

### v1

@@ II-B4 Software
Flux Framework is our choice of workload manager, as it is developed as the primary workload manager at Lawrence Livermore National Laboratory and will be provisioned to the upcoming El Capitan system in late 2024 [74]. Flux features graph-based scheduling and hierarchical workload management, which enables features like launching nested instances, and as previously mentioned, handles job submissions of up 100 million jobs [39, 52, 53]. It provides several key features that make it well-suited for this particular environment. In particular, having a shared filesystem is an extra setup step and incurred cost in a cloud environment, and since the workloads in this experiment did not require one, we decided to omit it. However, this choice made tasks such as sharing a file-based key from the control plane to other workers challenging. By using “flux archive” [75] to save the key to a memory map and “flux exec” [76] to send it to workers, the problem is solved, and a shared filesystem is not needed to share files, or even to run commands across workers. This second ability to execute commands across the cluster from a single node was required to start and connect each node in the Usernetes cluster. In early prototypes, each virtual machine needed to be connected to via a secure shell (ssh) to manually run the command. Now, with much larger clusters, the worker nodes never have to be accessed, as all interactions with them occur from the lead broker or control plane. One final bug with the Flux setup was the use of systemd. While the HPC family of AWS instances is explicitly intended for HPC use-cases and the memory limit is set to “unlimited” systemd by default will set a ceiling, and an additional parameter “LimitMEMLOCK=infinity” is required in the service file for Flux to also allow for unlimited access [77].
@@
The core of the setup requires a base of virtual machines, each of which is provisioned with an HPC workload manager, Flux Framework, the system requirements for Usernetes (Section II-A2), a clone of Usernetes that is ready for starting the cluster nodes, along with applications of interest. This setup affords three different models for running jobs across nodes – without Usernetes on “bare metal,” with Usernetes across pods and nodes, and with Usernetes, still across pods and nodes, but deploying the entirety of Flux Framework inside of the set of pods to afford the same hierarchical management. We put bare metal in quotes because a virtual machine is not explicitly bare metal, but the closest to it that can be achieved in cloud. For the first model, an example would be running an HPC MPI application directly with Flux Framework. For the second second model, we might use Flux to create a partition of nodes for Usernetes, and then run a multi-pod service across them such as a database or task queue. For the third model we create an equivalent partition of nodes, but then deploy the Flux Operator [36], a Kubernetes operator that deploys the entirety of Flux Framework inside of Kubernetes, within. As it sounds, this is running an HPC application inside of Flux, inside of Usernetes, inside of Flux. This final approach would require the HPC application to be containerized, and the Flux Operator adds the entire Flux install to it on the fly. For these experiments, we took care to ensure that the Flux inside of the container matched the outer install on the virtual machines exactly. This extra step was to ensure that the running of applications of interest using Flux in Usernetes was as similar as possible to the running of the same applications on the virtual machine directly.
@@
For networking, although the Flux Operator uses a headless Service [78], the ability of that service to be performant is dependent on the underlying network provided by Usernetes. For the Flux Operator to take full advantage of the Elastic Fiber Adapter, a DaemonSet [79] needed to be deployed that would expose the host EFA drivers to containers that were labeled to require the “vpc.amazonaws.com/efa” label under resource requests and limits [80]. Although this DaemonSet is provided by AWS to deploy EFA to traditional Kubernetes clusters, by way of not being directly on their instance types, we created a custom deployment configuration that removed the selectors and would thus deploy the DaemonSet to our Usernetes nodes.
@@
Figure 2: Networking Setup comparison between traditional user-space Kubernetes with and without the Elastic Fiber Adapter (EFA). For using Usernetes out of the box, a traditional path that uses TCP/IP is used, taking advantage of components in the blue, dotted box to afford rootless networking. This approach takes a performance hit because of the need to receive and process packets in the TAP device. The details of this approach are explained in Figure 1 of [44] from which the architecture is derived. By way of using EFA (bottom of diagram) applications use the user-space libfabric API to interact directly with hardware. This means that HPC applications do not take a hit as a result of the setup.
@@ II-C Experiments
The following experiments have two primary goals – to assess performance issues that might arise with running traditional HPC applications in Usernetes (Section II-C1) and to demonstrate a benefit that the setup allows – a hybrid environment that combines an HPC workload manager and user-space Kubernetes to run HPC workloads alongside services (Section II-C2).
@@ II-C1 Scaling of HPC Applications
We are first interested in testing the strong scaling of an HPC application, namely one with MPI, across cluster sizes of 32, 16, 8, and 4. While it would not be needed to run an MPI application in Usernetes given the availability of “bare metal” virtual machines, the goal of this experiment is to identify performance issues that can impact more realistic use cases such as model-parallel machine learning training or other programming models that will exercise the HPC network.
We created 33 EC2 virtual machines (Section II-B3) due to the need for Usernetes to have a control plane, which typically is intended to act as manager and explicitly labeled to not accept work [81]. For our HPC application experiment we first chose LAMMPS, the Large-scale Atomic/Molecular Massively Parallel Simulator [82] as a proxy application to demonstrate the ability of the setup to strong scale with MPI, and in particular for an application that requires AllReduce collective and send and receive paired calls [6]. A problem size of 16 x 16 x 8 was carefully chosen to account for the iterations, different scales, and final time and cost for the experiment runs. For a dedicated networking benchmark, we chose to use the OSU benchmarks [83] to measure MPI collective performance, bandwidth, and point to point latency. We decided to test each of these application on “bare metal” with the HPC workload manager Flux Framework to compare with an equivalent run using the Flux Operator in Usernetes. For the bare metal cases, we also chose to run a containerized variant using Singularity [18] as a point of comparison to provide evidence for or against containerization. This meant 20 iterations for each of the following cases, each being run with Flux, and across sizes 32, 16, 8, and 4:
@@ II-C2 Hybrid Demonstration of Environment Capabilities
The compelling use case for a hybrid environment with an HPC workload manager and user-space Kubernetes is being able to run HPC workloads alongside services. To best understand this approach, we will first describe a simple setup for running a simple, multi-node application in each space (Figure 3): Kubernetes, native HPC, and converged. In the Kubernetes setup, we use a “Job” abstraction with a headless Service to give a shared network to a set of Pods, each running a container with the application inside. This approach is advantageous for the orchestrated automation and modularity of applications and services, but is designed for loosely coupled, stateful services and not well suited to low latency, tightly coupled applications. It would be well-suited, however, for associated services like databases, task queues, and model APIs. The scheduler also is designed to support single Pods, and requires extensive modification to handle logical groups of work. Storage needs must use the container storage interface (CSI) drivers coupled with cloud solutions, and is challenging. If applications run, there is a loss of performance, and Kubernetes is not yet suited to handle fine-grained topology mapping that might be needed. For traditional HPC, the user submits jobs to the workload manager by way of a headless login node, and an allocation of resources is scheduled for the work. When the allocation is granted, the nodes are connected via a high speed interconnect networking fabric, and the application, which is typically built on these same machines, is run on the bare metal machines that are physically connected, and with a resource mapping that optimizes performance. Unlike Kubernetes, support for orchestration and automation is poor, with scientists often relying on workflow tools that use a shared filesystem to represent state, or bash scripts that utilize polling and custom checks. If a service such as a database or task queue is needed, it typically is a special request, as shared, open ports are not common or considered good practice for headless, private systems. However, this setup affords scalability, fine-grained resource mapping, and performance, which often are top priority for scientific simulations. Finally, a converged setup with a workload manager scheduling work to both bare metal HPC and user-space Kubernetes demonstrates the strength of combining the two. In this setup, user-space Kubernetes and traditional HPC simulations are running under the same allocation, and both are provisioned via a single batch job that uses a hierarchy to create sub-allocations for each. For this hypothetical example (Figure 3), two of the four nodes are being used for a service that is accessible to the simulations, and the simulations run alongside the service, using it. The batch job handles the creation and cleanup of associated services along with submitting jobs for the simulations. The three setups are shown side by side in Figure 3, and the advantages and disadvantages to each are shown in Table I.
@@
Figure 3: Running a simple, multi-node job across three setups: Kubernetes, traditional HPC, and a converged computing approach (user-space Kubernetes deployed as part of a batch job). Kubernetes (left panel) affords portability, orchestration, and modularity of services and applications at the cost of performance, reliable access to a file system, and fine-grained resource mapping. Traditional HPC workload managers (middle panel) are oriented toward applications that warrant tight coupling and performance, but this comes at the cost of portability, and easy orchestration of associated services. A converged approach with user-space Kubernetes that is deployed as part of a batch job offers the best of both worlds: bare metal HPC for performant, scaled simulation, and the modular, automated service architecture afforded by Kubernetes. Missing from this picture is the workflow tool or orchestration and scheduling of those two components.
As a multi-node job is too simple an example, we aimed to create a prototype setup with machine learning to fully demonstrate the advantages of the converged approach. Toward this goal, we developed a simple workflow that runs the same HPC application, LAMMPS, on bare metal with the workload manager Flux Framework, to predict LAMMPS wall times from the problem sizes. While the predictions themselves are not a contribution offered by this work, they are a key component of a simulation of the type of converged workload that we want to support and champion. As this setup aims to demonstrate the hybrid model, the specific problem size and number of nodes is less important than in our previous scaling study, and so we chose a cost effective cluster size of 5 nodes (allowing for LAMMPS to run on 4 nodes) and a problem size range between 1 and 8 for each of the dimensions for x, y, and z that go into LAMMPS. For this setup, LAMMPS jobs are submitted to the workload manager and on completion, the result data (the Y in the model being the LAMMPS wall times, and X features as the randomly chosen x, y, and z parameters that define a bounding box or problem size) to a machine learning service being run by Usernetes. This setup is depicted in Figure 4. Specifically, this figure would be a more detailed manifestation of the third panel in Figure 3.
@@
In practice, the hypothetical user simply needs to start the machine learning service, a Kubernetes Deployment [90], and then ensure that the port for it is exposed via Ingress [91], or directly from the port of the Pod mapped to the host. A simple script, ideally in a container for reproducibility, can then be called against the API endpoint to create the models, which in this case were three regression models. Once the models are created, another script (also ideally from a container) can submit one or more jobs directly to the workload manager that will run the HPC workload (LAMMPS) and then submit the resulting wall time (Y) and features (x, y, z) to the server training endpoint. The model trains, one point at a time, in this fashion, up until the user decides it has reached some metric of goodness or number of samples (also provided by the API server). When these submissions are done, a second batch is sent off to again run LAMMPS, but the actual wall times, Y, are left out, and the server is asked for a prediction based on the feature vector with the problem size, x, y, z. Since these jobs are actually running LAMMPS, it means that we have the true value of the wall time, and can use that to plot the actual vs. predicted times for each model. This final step can show how well the model did with an R squared value. Note that although simple regression models were chosen for this demonstration, online machine learning offers many different kinds of models. For our prototype example, we did 1000 training runs and 250 test cases, each randomly selecting problem sizes for x, y, z between 1 and 8. This range was chosen to pin the maximum runtime (and thus cloud cost) to a maximum value. A more rigorous (and expensive) study would warrant analyzing runtime characters of a larger range of problem sizes.
@@ III-A LAMMPS
We ran LAMMPS at a problem size of 16 x 16 x 8 across job sizes of 4, 8, 16, and 32 nodes, 20 iterations each, corresponding to 64, 128, 256, and 512 MPI ranks, respectively, and across 5 setups. This produced a total of 400 individual runs. We found no discernable differences between LAMMPS on bare metal, with or without Usernetes running, and running from within a container or directly on the virtual machine (Figure 5). Running the application in Usernetes produced runtimes that were only marginally slower across problem sizes, and with lower variability. Strong scaling produced a nice result, as the overall wall time decreased as the problem size was held constant and resources available increased. LAMMPS wall times, including means and standard deviations for each scoped experiment, are included in Table II.
Figure 5: LAMMPS runs with strong scaling for a 16 x 16 x 8 problem size and 20 iterations per experiment and size. There were no discernable differences between running LAMMPS on bare metal, with or without a container, and with or without Usernetes in the background, so these results are combined into a single group “hpc” (orange). LAMMPS running in Usernetes (blue) was only marginally slower with lower variability. Node sizes of 4, 8, 16, and 32 nodes correspond to MPI ranks 64, 128, 256, and 512 MPI, respectively. Means and standard deviations for each experiment type are included in Table II
@@
Nodes Ranks Time Seconds % CPU
@@ III-B OSU Benchmarks
We ran a subset of OSU Benchmarks across job sizes of 4, 8, 16, and 32 nodes, 20 iterations each, corresponding to 64, 128, 256, and 512 MPI ranks, respectively, and across 5 setups for each of AllReduce, Barrier, and point to point latency and bandwidth. Point to point bandwidth (osu_bw) was approximately 22.3% lower (meaning less data transferred) in Usernetes as compared to “bare metal” for the smallest message size of 1 byte (1.712 MB/s vs. 1.3 MB/s, respectively), and 0.32% lower for the largest message size of 4,194,304 bytes (approximately 24,125 MB/s vs. 24,202 MB/s, respectively) (Figure 6). For OSU Latency (osu_latency) we found that average latency for Usernetes is 65% greater for the smallest size (Figure 9) (7.46 and 12.31 microseconds, respectively). For collective call OSU Barrier (osu_barrier) Usernetes average latency is 31.89 microseconds higher for 4 nodes (Figure 7), which means the average latency for Usernetes is 78.68% greater. This particular benchmark had huge outliers across experiments, suggesting a common weakness that might relate to the physical location of one or more nodes. Across OSU benchmarks, bare metal performance was equivalent to performance running in a container, and the addition of Usernetes in the background added overhead that impacted the benchmarks only for collective benchmarks at larger sizes. This suggests that an allocation running Usernetes alongside bare metal HPC does incur a cost to the network, and should be considered with the setup. For the AllReduce benchmark (Figure 8) Usernetes average latency for 4 nodes was 3.6x to 13.7x higher, and for 32 nodes it was 2.89x to 4.32x higher (Figure 8).
Figure 6: The OSU Bandwidth point to point benchmark. Point to point bandwidth was approximately 22.3% lower (lower is less performant) in Usernetes as compared to bare metal for the smallest message size of 1 byte, and trivially 0.32% lower for the largest message size.
@@
Figure 9: The OSU Latency point to point benchmark had comparable performance across “bare-metal” setups, and 65% greater for the smallest size in Usernetes. The latency values overall, however, are relatively good for the cloud for all setups – with “bare-metal” (with and without container) experiments having an approximate average latency of 7.5 microseconds, and Usernetes 12.3 microseconds.

### v2

@@ II-B4 Software
We use the Flux Framework, a modular, hierarchical resource management and graph-based scheduling framework developed at the Lawrence Livermore National Laboratory. Flux supports scheduling dynamic resources via its graph-based representation [52], and its modular architecture facilitates integration with cloud [11, 14]. Its hierarchical resource management enables launching nested instances, which allows for high throughput and submission of over 100 million jobs [19, 32, 33]. Flux is scalable and designed to manage and schedule very large systems, such as the upcoming exascale El Capitan system [53].
@@
It provides several key features that make it well-suited for combination with Usernetes. In particular, having a shared filesystem is an extra setup step and incurred cost in a cloud environment, and since the workloads (Section 2) in this study did not require one, we decided to omit it. However, this choice made tasks such as sharing a file-based key from the control plane to other workers challenging. By using “flux archive” [54] to save the key to a memory map and “flux exec” [55] to send it to workers, the problem is solved, and a shared filesystem is not needed to share files, or even to run commands across workers. This second ability to execute commands across the cluster from a single node was required to start and connect each node in the Usernetes cluster. In early prototypes, each virtual machine needed to be connected to via a secure shell (ssh) to manually run the command. Now, with much larger clusters, the worker nodes never have to be accessed, as all interactions with them occur from the lead broker or control plane. One final bug with the Flux setup was the use of systemd. While the HPC family of AWS instances is explicitly intended for HPC use-cases and the memory limit is set to “unlimited” systemd by default will set a ceiling, and an additional parameter “LimitMEMLOCK=infinity” is required in the service file for Flux to also allow for unlimited access [56].
@@
The core of the setup requires a base of virtual machines, each of which is provisioned with an HPC workload manager, Flux Framework, the system requirements for Usernetes (Section II-A2), a clone of Usernetes that is ready for starting the cluster nodes, along with applications of interest. This setup affords three different models for running jobs across nodes – without Usernetes on “bare metal,” with Usernetes across pods and nodes, and with Usernetes, still across pods and nodes, but deploying the entirety of Flux Framework inside of the set of pods to afford the same hierarchical management. We put bare metal in quotes because a virtual machine is not truely bare metal, but the closest to it that can be achieved in cloud. For the first model, an example would be running an HPC MPI application directly with Flux Framework. For the second second model, we might use Flux to create a partition of nodes for Usernetes, and then run a multi-pod service across them such as a database or task queue. For the third model we create an equivalent partition of nodes, but then deploy the Flux Operator [14], a Kubernetes operator that deploys the entirety of Flux Framework inside of Kubernetes, within. As it sounds, this is running an HPC application inside of Flux, inside of Usernetes, inside of Flux. This final approach would require the HPC application to be containerized, and the Flux Operator adds the entire Flux install to it on the fly. For these experiments, we took care to ensure that the Flux inside of the container matched the outer install on the virtual machines. We ensured matching versions so the applications’ execution with Flux in Usernetes was as similar as possible to the same applications’ execution on the virtual machine.
@@
For networking, although the Flux Operator uses a headless Service [57], its performance depends on the underlying network provided by Usernetes. For the Flux Operator to take full advantage of the Elastic Fabric Adapter, a DaemonSet [58] needed to be deployed that would expose the host EFA drivers to containers that were labeled to require the “vpc.amazonaws.com/efa” label under resource requests and limits [59]. Although this DaemonSet is provided by AWS to deploy EFA to traditional Kubernetes clusters, by way of not being directly on their instance types, we created a custom deployment configuration that removed the selectors and would thus deploy the DaemonSet to our Usernetes nodes.
@@
Figure 2: Networking Setup comparison between traditional user-space Kubernetes with and without the Elastic Fabric Adapter (EFA). For using Usernetes out of the box, a traditional path that uses TCP/IP is used, taking advantage of components in the blue, dotted box to afford rootless networking. This approach takes a performance hit because of the need to receive and process packets in the TAP device. The details of this approach are explained in Figure 1 of [24] from which the architecture is derived. By way of using EFA (bottom of diagram) applications use the user-space libfabric API to interact directly with hardware. This means that HPC applications do not take a hit as a result of the setup.
@@ II-C Experiments
The following experiments have two primary goals – to assess performance issues that might arise with running traditional HPC applications in Usernetes (Section II-C1) and to demonstrate a benefit that the setup affords – a hybrid environment that combines an HPC workload manager and user-space Kubernetes to run HPC workloads alongside services (Section II-C2).
@@ II-C1 Scaling of HPC Applications
We are first interested in testing the strong scaling of an HPC application, namely one with MPI, across cluster sizes of 4, 8, 16, and 32. While running an MPI application in Usernetes is not strictly necessary given the availability of “bare metal” virtual machines, the goal of this experiment is to identify performance issues that can impact more realistic use cases such as model-parallel machine learning training or other programming models that will exercise the HPC network.
We created 33 EC2 virtual machines (Section II-B3) due to the need for Usernetes to have a control plane, which typically is intended to act as manager and explicitly labeled to not accept work [60]. For our HPC application experiment we first chose LAMMPS, the Large-scale Atomic/Molecular Massively Parallel Simulator [61] as a proxy application to demonstrate the ability of the setup to strong scale with MPI, and in particular for an application that requires AllReduce collective and send and receive paired calls [62]. A problem size of $16\times 16\times 8$ was chosen to account for the iterations, different scales, and final time and cost for the experiment runs. For a dedicated networking benchmark, we chose to use the OSU benchmarks [63] to measure MPI collective performance, bandwidth, and point to point latency. We decided to test each of these application on “bare metal” with the HPC workload manager Flux Framework to compare with an equivalent run using the Flux Operator in Usernetes. We also chose to run a containerized variant on bare metal using Singularity [64] to provide a point of performance comparison. We ran 20 iterations with Flux across sizes 4, 8, 16, and 32 for each of the following cases:
@@ II-C2 Hybrid Demonstration of Environment Capabilities
The compelling use case for a hybrid environment with an HPC workload manager and user-space Kubernetes is being able to run HPC workloads alongside services within a user-scoped allocation. To best understand this approach, we will first describe a simple setup for running a simple, multi-node application in each space (Figure 3): Kubernetes, native HPC, and converged. In the Kubernetes setup, we use a “Job” abstraction with a headless Service to give a shared network to a set of Pods, each running a container with the application inside. This approach is advantageous for the orchestrated automation and modularity of applications and services, but is designed for loosely coupled, stateful services and may not be best suited to low latency, tightly coupled applications. It would be well-suited, however, for associated services like databases, task queues, model APIs, and AI/ML components of a scientific workflow. The scheduler also is designed to support single Pods, and requires extensive modification to handle logical groups of work. Storage needs must use the container storage interface (CSI) drivers coupled with cloud solutions, and are challenging. If applications run, there is a loss of performance, and Kubernetes is not yet suited to handle fine-grained topology mapping that might be needed.
For traditional HPC, the user submits jobs to the workload manager by way of a headless login node, and an allocation of resources is scheduled for the work. When the allocation is granted, the nodes are connected via a high speed interconnect networking fabric, and the application, which is typically built on these same machines, is run on the bare metal machines that are physically connected, and with a resource mapping that optimizes performance. Unlike Kubernetes, support for orchestration and automation is poor, with scientists often relying on workflow tools that use a shared filesystem to represent state, or bash scripts that utilize polling and custom checks. If a service such as a database or task queue is needed, it typically is a special request, as shared, open ports are not common or considered good practice for headless, private systems. However, this setup affords scalability, fine-grained resource mapping, and performance, which often are top priority for scientific simulations.
Finally, a converged setup with a workload manager scheduling work to both bare metal HPC and user-space Kubernetes demonstrates the strength of combining the two. In this setup, user-space Kubernetes and traditional HPC simulations are running under the same allocation, and both are provisioned via a single batch job that uses a hierarchy to create sub-allocations for each. For this hypothetical example (Figure 3), two of the four nodes are being used for a service that is accessible to the simulations, and the simulations run alongside the service, using it. The batch job handles the creation and cleanup of associated services along with submitting jobs for the simulations. The three setups are shown side by side in Figure 3, and the advantages and disadvantages to each are shown in Table I.
@@
Figure 3: Running a simple, multi-node job across three setups: Kubernetes, traditional HPC, and a converged computing approach (user-space Kubernetes deployed as part of a batch job). Kubernetes (left panel) affords portability, orchestration, and modularity of services and applications at the cost of performance, reliable access to a file system, and fine-grained resource mapping. Traditional HPC workload managers (middle panel) are oriented toward applications that require tight coupling and performance, which comes at the cost of portability and easy orchestration of associated services. A converged approach with user-space Kubernetes that is deployed as part of a batch job offers the best of both worlds: bare metal HPC for performant, scaled simulation, and the modular, automated service architecture afforded by Kubernetes. Missing from this picture is the workflow tool or orchestration and scheduling of those two components.
As a multi-node job is too simple an example, we aimed to create a prototype setup with machine learning to fully demonstrate the advantages of the converged approach. Toward this goal, we developed a simple workflow that runs the same HPC application, LAMMPS, on bare metal with the workload manager Flux Framework, to predict LAMMPS wall times from the problem sizes. While the predictions themselves are not a contribution offered by this work, they are a key component of a proxy to the type of converged workload that we want to support and champion. As this setup aims to demonstrate the hybrid model, the specific problem size and number of nodes is less important than in our previous scaling study, and so we chose a cost effective cluster size of 5 nodes (allowing for LAMMPS to run on 4 nodes) and a problem size range between 1 and 8 for each of the dimensions for x, y, and z that go into LAMMPS. For this setup, LAMMPS jobs are submitted to the workload manager and on completion, the result data (the Y in the model being the LAMMPS wall times, and X features as the randomly chosen x, y, and z parameters that define a bounding box or problem size) to a machine learning service being run by Usernetes. This setup is depicted in Figure 4. Specifically, this figure would be a more detailed manifestation of the third panel in Figure 3.
@@
In practice, the hypothetical user simply needs to start the machine learning service, a Kubernetes Deployment [71], and then ensure that the port for it is exposed via Ingress [72], or directly from the port of the Pod mapped to the host. A simple script, ideally in a container for reproducibility, can then be called against the API endpoint to create the models, which in this case were three regression models. Once the models are created, another script (also ideally from a container) can submit one or more jobs directly to the workload manager that will run the HPC workload (LAMMPS) and then submit the resulting wall time (Y) and features (x, y, z) to the server training endpoint. The model trains, one point at a time, in this fashion, up until the user decides it has reached some metric of goodness or number of samples (also provided by the API server). When these submissions are done, a second batch is sent off to again run LAMMPS, but the actual wall times, Y, are left out, and the server is asked for a prediction based on the feature vector with the problem size, x, y, z. Since these jobs are actually running LAMMPS, it means that we have the true value of the wall time, and can use that to plot the actual vs. predicted times for each model. This final step can show how well the model did with an $R^{2}$ value. Note that although simple regression models were chosen for this demonstration, online machine learning offers many different kinds of models. For our prototype example, we did 1000 training runs and 250 test cases, each randomly selecting problem sizes for x, y, z between 1 and 8. This range was chosen to pin the maximum runtime (and thus cloud cost) to a maximum value. A more rigorous (and expensive) study would warrant analyzing runtime characters of a larger range of problem sizes.
@@ III-A LAMMPS
We ran LAMMPS at a problem size of $16\times 16\times 8$ across job sizes of 4, 8, 16, and 32 nodes, 20 iterations each, corresponding to 64, 128, 256, and 512 MPI ranks, respectively, and across 5 setups. This produced a total of 400 individual runs. LAMMPS running on bare metal is compared to LAMMPS running in Usernetes in 5. LAMMPS on bare metal was similar with or without Usernetes running, and running from within a container or directly on the virtual machine. We observed LAMMPS to strong scale; LAMMPS wall times, including means and standard deviations for each scoped experiment are included in Table II.
Figure 5: LAMMPS runs with strong scaling for a $16\times 16\times 8$ problem size and 20 iterations per experiment and size. LAMMPS running in Usernetes was only marginally slower with lower variability. Node sizes of 4, 8, 16, and 32 nodes correspond to MPI ranks 64, 128, 256, and 512 MPI, respectively. Containerized experiments and those with Usernetes running in the background yielded similar results. Means and standard deviations for all experiments are included in Table II
@@
Nodes Ranks Mean (s) (stddev) % CPU
@@ III-B OSU Benchmarks
We ran a subset of OSU Benchmarks across job sizes of 4, 8, 16, and 32 nodes, 20 iterations each, corresponding to 64, 128, 256, and 512 MPI ranks, respectively, and across 5 setups for each of AllReduce, Barrier, and point to point latency and bandwidth. Point to point bandwidth (osu_bw) was approximately 22.3% lower (meaning less data transferred) in Usernetes as compared to “bare metal” for the smallest message size of 1 byte (1.712 MB/s vs. 1.3 MB/s, respectively), and 0.32% lower for the largest message size of 4,194,304 bytes (approximately 24,125 MB/s vs. 24,202 MB/s, respectively) (Figure 6). For OSU Latency (osu_latency) we found that average latency for Usernetes is 65% greater for the smallest size (Figure 9) (7.46 and 12.31 microseconds, respectively). For collective call OSU Barrier (osu_barrier) Usernetes average latency is 31.89 microseconds higher for 4 nodes (Figure 7), which means the average latency for Usernetes is 78.68% greater. This particular benchmark had huge outliers across experiments, suggesting a common weakness that might relate to the physical location of one or more nodes. Across OSU benchmarks, bare metal performance was equivalent to performance running in a container, and the addition of Usernetes in the background added overhead that impacted the benchmarks only for collective benchmarks at larger sizes. This suggests that an allocation running Usernetes alongside bare metal HPC does incur a cost to the network, and should be considered with the setup. For the AllReduce benchmark (Figure 8) Usernetes average latency for 4 nodes was 3.6×3.6$\times$3.6 × to 13.7×13.7$\times$13.7 × higher, and for 32 nodes it was 2.89×2.89$\times$2.89 × to 4.32×4.32$\times$4.32 × higher (Figure 8).
Figure 6: The OSU Bandwidth point to point benchmark. Point to point bandwidth was approximately 22.3% lower (lower is less performant) in Usernetes as compared to bare metal for the smallest message size of 1 byte, and 0.32% lower for the largest message size.
@@
Figure 9: The OSU Latency point to point benchmark had comparable performance across “bare-metal” setups, and 65% greater for the smallest size in Usernetes. The “bare-metal” (with and without container) experiments have an average latency of 7.5 microseconds, and Usernetes 12.3 microseconds.